OpenStudy (anonymous):

how do I write a equation in standard form for "passes through -2 -1 and 6 -4

OpenStudy (anonymous):

are those pairs?

OpenStudy (anonymous):

yeah bo they are (-2 ,-1) (6 ,-4)

OpenStudy (anonymous):

standard form is Ax+By=C, I like to start with the linear form y=mx+b First we the slope equations (-4+1) /(6+2)=-3/8 so now we have y=-3/8+b we just plug in one of the coords, lets do (-2,-1) -1=(-3/8)x-2 +b -1=(-3/4)+b add 3/4 to both sides b=1/4 now we have y=(-3/8)x+1/4 now we just rearrange to standard Ax+By=C C is 1/4 add 3/8x to both sides y+(3/8)x=1/4 now my teachers usually dont like fractions in standard form so lets multiply by 8 8y+3x=2 Ta da!